OpenStudy (dls):

A body falls from rest.It covers as much distance in the last second of its motion as covered d in the first 3 seconds.The body has fallen for a time of___

OpenStudy (dls):

Is it 3 seconds?

OpenStudy (dls):

@Yahoo!

OpenStudy (anonymous):

Do u knw the formula for the dist travelled in nth second

OpenStudy (dls):

Yes

OpenStudy (anonymous):

can u write that

OpenStudy (dls):

S=u+1/2(2n-1)

OpenStudy (anonymous):

u certainly misses out "a"

OpenStudy (dls):

sorry yeah..

OpenStudy (dls):

just confirm my answer :(

OpenStudy (anonymous):

wat is ur answer

OpenStudy (dls):

3?

OpenStudy (anonymous):

i dont think it is 3

OpenStudy (anonymous):

@experimentX @mukushla @ajprincess

OpenStudy (anonymous):

s1 = 1/2 * a * t^2

OpenStudy (anonymous):

= 5 * 9 = 45 m

OpenStudy (anonymous):

dn = u + a/2 [ 2n-1]

OpenStudy (anonymous):

45 = 5 * { 2n -1) 9 = 2n -1 n = 5

OpenStudy (anonymous):

so..5 seconds

OpenStudy (dls):

shouldn't it be 2(n-1-1) ? how do we know nth second means last second here?

OpenStudy (anonymous):

did nt understand ur question lol

OpenStudy (anonymous):

Formula for the distance travelled in n th second where n = 1 ,2 , 3....................................dn = u + a/2 [ 2n-1]

OpenStudy (anonymous):

nop @ajprincess dn = u + a/2 [ 2n-1] = dn = u + a [ n-1/2]

OpenStudy (ajprincess):

ya u right.@yahoo!. Sorry abt the interruption. carry on.

OpenStudy (anonymous):

no problem

OpenStudy (dls):

what according to u is n in 2n-1 ?

OpenStudy (anonymous):

n === time

OpenStudy (dls):

n means nth second.............

OpenStudy (anonymous):

yup.....

OpenStudy (dls):

but here we are asked n-1 second so.. shouldn't it be 2n-1-1?

OpenStudy (anonymous):

so...5 th second is last second

OpenStudy (dls):

im confused nvm :/

OpenStudy (anonymous):

just read the question

OpenStudy (anonymous):

It covers as much distance in the last second of its motion as covered d in the first 3 seconds

OpenStudy (dls):

thanks

OpenStudy (anonymous):

"last second of its motion"

OpenStudy (anonymous):

u understood??